# Efficient Reasoning Reuse, Don’t Recompute: Efficient Large Reasoning Model Inference via Memory Orchestration

### v1

@@ 1 Introduction
Large reasoning models (LRMs) have made “thinking longer” a default recipe for better answers: expand the chain-of-thought (CoT) [Wei2022CoT], sample more drafts [Wang2023SelfConsistency], and hope consensus emerges. Although this principle yields in real gains, it doesn’t come without costs. Lengthy test-time reasoning inflates latency, burns tokens on evidence the model already “knows,” and turns deployment into an exercise in budget triage rather than engineering. If LRMs are to be used in constrained settings, we need a way to spend fewer tokens without compromising accuracy.
@@ 2 Related Work
Research on long-term memory for language models spans non-parametric retrieval, structured graphs, and system-level abstractions. Early work coupled frozen models with dense or lexical retrieval [Khandelwal2020, Lewis2020, Guu2020], improving factual recall but often relying on heuristic chunking and calibration. Retrieval-pretrained and nearest-neighbor approaches scale this line further [Borgeaud2022], trading simplicity for freshness and editability. Structured methods instead frame memory as a graph, capturing entities and relations to support multi-hop reasoning [Anokhin2024, Li2024GraphReader], while system-level proposals treat memory as a schedulable resource with paging and lifecycle management [Packer2023, Li2025MemOS]. These designs extend capacity but typically return verbose snippets, leaving models to repeatedly re-narrate evidence.
Parallel to advances in memory, work on LRMs has begun to emphasize efficiency. While early advances in LRMs emphasized thinking longer through extended chains or ensemble-style decoding, more recent work has shifted toward making inference more efficient. Methods such as Skeleton-of-Thought [Ning2023SoT] and Chain-of-Draft [Xu2025CoD] restructure reasoning to reduce redundancy and accelerate generation, showing that careful control over inference can yield accuracy at lower cost. Our approach is complementary: rather than restructuring the chain itself, we compress the evidence surface, reducing the need for redundant reasoning by constraining what models are allowed to see and cite.
@@ 3 The ENGRAM-R Architecture
We extend the ENGRAM memory architecture [Anonymous2026] into a full efficiency layer for LRMs. Analogous to ENGRAM, our system organizes dialogue into typed memory stores [Tulving1972, Cohen1980] and performs dense retrieval at query time, but we add two critical extensions: (i) retrieved items are re-rendered into compact Fact Cards, and (ii) models are instructed to cite cards directly in their reasoning trace. These additions turn memory from a verbose context dump into a bounded, auditable evidence substrate. The pipeline consists of five main stages, marked (1)–(5) in Figure 1 and are referenced throughout this section. Here, we briefly recap the essentials and then introduce our extensions.
@@ 3.1 Problem Setup
A naive baseline is to pass the entire conversation as input. This guarantees recall but scales poorly: tokens grow linearly with turns, often exceeding context limits. Test-time scaling (TTS) strategies such as extended chains [Wei2022CoT] or self-consistency [Wang2023SelfConsistency] amplify cost further, inflating both latency and token budget. To avoid this, ENGRAM uses retrieval:
@@ 4.1 Benchmarks
LoCoMo. LoCoMo consists of multi-session dialogues built through a human–machine pipeline grounded in personas and event graphs, then refined by human annotators for long-range consistency [Maharana2024]. It contains 10 dialogues, each averaging 600 turns and 16K tokens across up to 32 sessions. The QA split covers five categories: single-hop, multi-hop, temporal, open-domain, and adversarial. In line with prior work, we exclude adversarial cases when reporting QA metrics and present category-level results for the remaining four.
LongMemEval. LongMemEval is designed to test interactive memory in user–assistant dialogues, covering five abilities: information extraction, multi-session reasoning, temporal reasoning, knowledge updates, and abstention (declining to answer when evidence is insufficient) [Wu2024LongMemEval]. It contains 500 curated questions embedded in chat histories of configurable length. For our study, we focus on LongMemEval (115K tokens per problem) and report QA metrics on the multi-session reasoning and temporal reasoning categories. These settings are particularly challenging and better highlight whether memory orchestration can deliver efficiency gains without sacrificing accuracy.
@@ 4.2 Backbone and baselines
All experiments are conducted with gpt-oss-20b [gptoss2025], selected as a strong open-source LRM with competitive chain-of-thought capabilities. The model weights remain frozen; our interventions act only at inference time. We evaluate two inference settings:
@@
ENGRAM-R. This setting augments inference with a compact memory orchestration layer: interaction traces and knowledge are organized into typed records; for a given query, we perform dense retrieval and truncate to a fixed evidence budget before presenting this condensed context to the model. The memory layer runs with gpt-4o-mini [OpenAI2023], a non-reasoning, intentionally fast and lightweight model, as its backbone, conserving reasoning for the question-answering stage.
@@ 4.3 Metrics
Quality. We use an LLM-as-Judge protocol [Zheng2023LLMJudge] as the main metric of semantic correctness. This involves using an independent LLM that, given the question, gold answer, and prediction, renders a binary semantic-correctness decision based on factual fidelity, completeness, and contextual appropriateness of the response. This ensures robustness to surface-level variation in responses.
@@ 5.5 Practical application
Beyond long-context stress tests, we evaluate a deployment-style scenario in healthcare, using the HealthBench benchmark [Arora2025HealthBench], to assess whether inference-time efficiency transfers to a high-stakes domain with shorter, multi-turn inputs. Even though the room to compress reasoning is smaller, ENGRAM-R consistently trims generative overhead while keeping accuracy near parity, with gains on temporally anchored questions. Full results and a separate discussion appear in Appendix C.
@@ A.5 Dense retrieval and budgeted aggregation (3–4)
Given a query , ENGRAM embeds it as and retrieves candidates within each store by cosine similarity [Karpukhin2020]
@@
In all experiments we default to , a knee point identified by a -sweep ablation performed in the ENGRAM paper [Anonymous2026].

### v2

@@ 1 Introduction
Large reasoning models (LRMs) have made “thinking longer” a default recipe for better answers: expand the chain-of-thought (CoT) [19], sample more drafts [18], and hope consensus emerges. Although this principle yields in real gains, it doesn’t come without costs. Lengthy test-time reasoning inflates latency, burns tokens on evidence the model already “knows,” and turns deployment into an exercise in budget triage rather than engineering. If LRMs are to be used in constrained settings, we need a way to spend fewer tokens without compromising accuracy.
@@ 2 Related Work
Research on long-term memory for language models spans non-parametric retrieval, structured graphs, and system-level abstractions. Early work coupled frozen models with dense or lexical retrieval [7, 8, 5], improving factual recall but often relying on heuristic chunking and calibration. Retrieval-pretrained and nearest-neighbor approaches scale this line further [3], trading simplicity for freshness and editability. Structured methods instead frame memory as a graph, capturing entities and relations to support multi-hop reasoning [1, 9], while system-level proposals treat memory as a schedulable resource with paging and lifecycle management [15, 10]. These designs extend capacity but typically return verbose snippets, leaving models to repeatedly re-narrate evidence.
Parallel to advances in memory, work on LRMs has begun to emphasize efficiency. While early advances in LRMs emphasized thinking longer through extended chains or ensemble-style decoding, more recent work has shifted toward making inference more efficient. Methods such as Skeleton-of-Thought [12] and Chain-of-Draft [21] restructure reasoning to reduce redundancy and accelerate generation, showing that careful control over inference can yield accuracy at lower cost. Our approach is complementary: rather than restructuring the chain itself, we compress the evidence surface, reducing the need for redundant reasoning by constraining what models are allowed to see and cite.
@@ 3 The ENGRAM-R Architecture
We extend the ENGRAM memory architecture [16] into a full efficiency layer for LRMs. Analogous to ENGRAM, our system organizes dialogue into typed memory stores [17, 4] and performs dense retrieval at query time, but we add two critical extensions: (i) retrieved items are re-rendered into compact Fact Cards, and (ii) models are instructed to cite cards directly in their reasoning trace. These additions turn memory from a verbose context dump into a bounded, auditable evidence substrate. The pipeline consists of five main stages, marked (1)–(5) in Figure 1 and are referenced throughout this section. Here, we briefly recap the essentials and then introduce our extensions.
@@ 3.1 Problem Setup
A naive baseline is to pass the entire conversation as input. This guarantees recall but scales poorly: tokens grow linearly with turns, often exceeding context limits. Test-time scaling (TTS) strategies such as extended chains [19] or self-consistency [18] amplify cost further, inflating both latency and token budget. To avoid this, ENGRAM uses retrieval:
@@ 4.1 Benchmarks
LoCoMo. LoCoMo consists of multi-session dialogues built through a human–machine pipeline grounded in personas and event graphs, then refined by human annotators for long-range consistency [11]. It contains 10 dialogues, each averaging 600 turns and 16K tokens across up to 32 sessions. The QA split covers five categories: single-hop, multi-hop, temporal, open-domain, and adversarial. In line with prior work, we exclude adversarial cases when reporting QA metrics and present category-level results for the remaining four.
LongMemEval. LongMemEval is designed to test interactive memory in user–assistant dialogues, covering five abilities: information extraction, multi-session reasoning, temporal reasoning, knowledge updates, and abstention (declining to answer when evidence is insufficient) [20]. It contains 500 curated questions embedded in chat histories of configurable length. For our study, we focus on LongMemEval (115K tokens per problem) and report QA metrics on the multi-session reasoning and temporal reasoning categories. These settings are particularly challenging and better highlight whether memory orchestration can deliver efficiency gains without sacrificing accuracy.
@@ 4.2 Backbone and baselines
All experiments are conducted with gpt-oss-20b [14], selected as a strong open-source LRM with competitive chain-of-thought capabilities. The model weights remain frozen; our interventions act only at inference time. We evaluate two inference settings:
@@
ENGRAM-R. This setting augments inference with a compact memory orchestration layer: interaction traces and knowledge are organized into typed records; for a given query, we perform dense retrieval and truncate to a fixed evidence budget before presenting this condensed context to the model. The memory layer runs with gpt-4o-mini [13], a non-reasoning, intentionally fast and lightweight model, as its backbone, conserving reasoning for the question-answering stage.
@@ 4.3 Metrics
Quality. We use an LLM-as-Judge protocol [22] as the main metric of semantic correctness. This involves using an independent LLM that, given the question, gold answer, and prediction, renders a binary semantic-correctness decision based on factual fidelity, completeness, and contextual appropriateness of the response. This ensures robustness to surface-level variation in responses.
@@ 5.5 Practical application
Beyond long-context stress tests, we evaluate a deployment-style scenario in healthcare, using the HealthBench benchmark [2], to assess whether inference-time efficiency transfers to a high-stakes domain with shorter, multi-turn inputs. Even though the room to compress reasoning is smaller, ENGRAM-R consistently trims generative overhead while keeping accuracy near parity, with gains on temporally anchored questions. Full results and a separate discussion appear in Appendix C.
@@ A.5 Dense retrieval and budgeted aggregation (3–4)
Given a query , ENGRAM embeds it as and retrieves candidates within each store by cosine similarity [6]
@@
In all experiments we default to , a knee point identified by a -sweep ablation performed in the ENGRAM paper [16].